# A conjecture on descents, inversions and the weak order on Coxeter groups

### v2

@@ 2.3. Bipartitions of the longest element
Recall that $W$ is finite if and only if there is an element of maximal length, which is unique and denoted by $w_{\circ}$. It is well-known that the longest element have the following properties: $\Phi^{+}=\Phi(w_{\circ})$; $w_{\circ}Sw_{\circ}=S$, $D_{R}(w_{\circ})=D_{L}(w_{\circ})=S$. Moreover, for all $w\in W$, we have:
@@
The next statement shows, in particular, that in the case of finite Weyl groups Conjecture 1 is equivalent to Ressayre’s statement ()($\triangle$)( △ ) from the introduction.
@@ 2.5. Conjecture 1, bipartitions and rectangled intervals in the weak order
The following statement shows that the maximal distance in an interval $[e,w]_{R}$ in $(W,\leq_{R})$ is achieved by a geodesic (a smallest path in the right Cayley graph) between $e$ and $w$.
@@
We believe that Proposition 2.12 is well-known, but we could not find a reference so we included a proof. More generally, it is known that the interval $[e,w]_{R}$ is convex, i.e., any geodesic between two elements of $[e,w]_{R}$ is contained in $[e,w]_{R}$. This follows from the interpretation of the inversion set $T(g)$, for $g\in W$, as the set $\mathcal{H}(g)$ of hyperplanes separating $g$ from the identity $e$. One can show that, for $u,v\in[e,w]_{R}$, the set of hyperplanes $\mathcal{H}(u,v)$ that separates $u$ from $v$ satisfies $d(u,v)=|\mathcal{H}(u,v)|$ and $\mathcal{H}(u,v)\subseteq\mathcal{H}(u)\cup\mathcal{H}(v)\subseteq\mathcal{H}(w)$. Moreover, any $g\in W$ in a geodesic from $u$ to $v$ satisfies $\mathcal{H}(u,g)\sqcup\mathcal{(}g,v)=\mathcal{H}(u,v)\subseteq\mathcal{H}(w)$ and $\mathcal{H}(g)\setminus\mathcal{H}(u,v)=H(u)\cup\mathcal{H}(v)\setminus\mathcal{H}(u,v)$. Therefore $\mathcal{H}(g)\subseteq\mathcal{H}(w)$, implying $g\leq_{R}w$.
Proof of Propostion 1.8.
Let $u,v,w\in W$. We need to show that $\{u,v\}$ is a bipartition of $w$ if and only if $\{u,v\}$ is a diameter of $[e,w]_{R}$.
Assume first that $\{u,v\}$ is a bipartition of $w$. Therefore $\ell(w)=\ell(u)+\ell(v)$ by Proposition 2.1 (2). Now, by Lemma 2.9, $\Phi(u)\cap\Phi(v)=\emptyset$. Finally we obtain $d(u,v)=\ell(u)+\ell(v)=\ell(w)$ and, therefore, $\{u,v\}$ is a diameter of $[e,w]_{R}$.
Assume now that $\{u,v\}$ is a diameter of $[e,w]_{R}$. On one hand, by definition and Lemma 2.11, we have:
@@
Together with ()($\star$)( ⋆ ), we obtain $-|\Phi(u)\cap\Phi(v)|\geq 0$. Therefore $\Phi(u)\cap\Phi(v)=\emptyset$. So by ()($\star$)( ⋆ ) again, we have $\ell(w)=|\Phi(w)|=|\Phi(u)\sqcup\Phi(v)|$. Since $\Phi(u)\cup\Phi(v)\subseteq\Phi(w)$, we conclude that $\Phi(w)=\Phi(u)\sqcup\Phi(v)$. In other words, $\{u,v\}$ is a bipartition of $w$. ∎
We end this section with the following proposition that proves the equivalence between Conjecture 2 and Conjecture 3.
@@
In particular, $\{u,v\}$ is a diameter of $[e,w]_{R}$ if and only if $\{(\varphi_{w}(u),\varphi_{w}(v)\}$ is a diameter of $[e,w^{-1}]_{R}$
@@ 3. The case of symmetric groups (Type A)
A permutation $\sigma\in\mathcal{S}_{n}$, as a bijection on the set $[n]=\{1,\dots,n\}$, is written as a word on the alphabet $[n]$: $\sigma=\sigma(1)\sigma(2)\dots\sigma(n)$ (the one-line notation). The (left) inversion set of $\sigma$ has therefore the following well-known interpretation:
@@
since in cycle notation $(i\ j)=(j\ i)$. The descent sets are $D_{L}(\sigma)=\{\tau_{3},\tau_{5},\tau_{7}\}$ and $D_{R}(\sigma)=\{\tau_{4},\tau_{5},\tau_{7}\}$. Therefore $\sigma$ has descents in positions $4,5$ and $7$.
@@ 3.1. Standardization of words, inversions and descents
(1) $\operatorname{invs}(w):=\{(w_{j},w_{i})\mid i<j\text{ and }w_{i}>w_{j}\}=\{(w_{j},w_{i})\mid w=\dots w_{i}\dots w_{j}\dots\}.$
@@
The word $w=391427$ has $2$ descents in positions $2$ and $4$ and $\operatorname{invs}(w)=\{(1,3),\ (1,9),\ (4,9),\ (2,3),\ (2,9),\ (2,4),\ (7,9)\}$.
@@
In the definition of inversions of the word $w$, we use the notation of a couple $(a,b)$ for an inversion instead of the cycle notation for which $(a\ b)=(b\ a)$. In particular, if $w\in\mathcal{S}_{n}$, then $T(w)=\{(a\ b)\mid(a,b)\in\operatorname{invs}(w)\}$.
The reason is that, in the proof of Theorem 1.5, we need to keep track of the following fact: $(a,b)\in\operatorname{invs}(w)$ means that the letter $a$ is strictly smaller that the letter $b$ and the letter $a$ appears after the letter $b$ in $w$, i.e., $w=\dots b\dots a\dots$.
@@
Let $I$ be a set of tuples $(a,b)$ with $1\leq a<b\leq n$ for some $n\geq 1$. Then $I$ is the inversion set of a certain permutation of $\mathcal{S}_{n}$ if and only if it satisfies the following for all $1\leq a<b<c\leq n$
@@
Let $w=526341$. It has $10$ inversions : $(1,2),(1,3),(1,4),(1,5),(1,6),(2,5),(3,5),(3,6),(4,5),(4,6)$ and $3$ descents in positions $1$, $3$, and $5$. Then $u=234561$ and $v=152634$ form a bipartition of $w$. Indeed, the inversions of $u$ are $(1,2),(1,3),(1,4),(1,5),(1,6)$ while the inversions of $v$ are $(2,5),(3,5),(3,6),(4,5),(4,6)$.
@@
Let $w$ be a permutation of $\mathcal{S}_{n}$ with $n\geq 1$. The decreased word of $w$ is obtained by removing the letter $n$ from $w$ and is denoted by $\tilde{w}$.
For an example, if $w=526341$, then $\tilde{w}=52341$.
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$ with $n\geq 1$, then $\{\tilde{u},\tilde{v}\}$ is a bipartition of $\tilde{w}$.
@@
Note that for $w\in S_{n}$ with $n\geq 1$, then $\operatorname{invs}(\tilde{w})=\{(a,b)\in\operatorname{invs}(w);b\neq n\}$. In other words, we remove the inversions containing $n$. In particular, as $\operatorname{invs}(w)=\operatorname{invs}(u)\sqcup\operatorname{invs}(v)$ we have $\operatorname{invs}(\tilde{w})=\operatorname{invs}(\tilde{u})\sqcup\operatorname{invs}(\tilde{v})$, which concludes the proof. ∎
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. The right letters of $w$ are defined as:
$\mathcal{R}(w):=\{a\mid(a,n)\in\operatorname{invs}(w)\}\cup\{n\}$
The left letters of $w$ are the complement in $[n]$ of the right letters:
$\mathcal{L}(w):=\{b\mid 1\leq b<n,(b,n)\not\in\operatorname{invs}(w)\}.$
@@
$w_{\mathcal{L}}$ $=\operatorname{std}(w_{|\mathcal{L}(w)}),$ $w_{\mathcal{R}}$ $=\operatorname{std}(w_{|\mathcal{R}(w)})$
$u_{\mathcal{L}}$ $=\operatorname{std}(u_{|\mathcal{L}(w)}),$ $u_{\mathcal{R}}$ $=\operatorname{std}(u_{|\mathcal{R}(w)})$
$v_{\mathcal{L}}$ $=\operatorname{std}(v_{|\mathcal{L}(w)}),$ $v_{\mathcal{R}}$ $=\operatorname{std}(v_{|\mathcal{R}(w)}).$
In other words, we cut $w$ into two parts (left and right) just before the maximal letter. Then we select those letters in $u$ and $v$ and standardize to form new permutations. Note that $n$ is by definition a right letter. The set $\mathcal{L}(w)$ of left letters might be empty, in this case the left words are the empty word and the right words are the original words $u,v,w$. If $n\geq 1$, the set $\mathcal{R}(w)$ of right letters is never the empty set as it contains at least the letter $n$. Note also that the left (resp. right) words of $u$ and $v$ depend on the left (resp. right) letters of $w$.
@@
$w_{\mathcal{L}}$ $=\operatorname{std}(52)=21,$ $w_{\mathcal{R}}$ $=\operatorname{std}(6341)=4231,$
$u_{\mathcal{L}}$ $=\operatorname{std}(25)=12,$ $u_{\mathcal{R}}$ $=\operatorname{std}(3461)=2341,$
$v_{\mathcal{L}}$ $=\operatorname{std}(52)=21,$ $v_{\mathcal{R}}$ $=\operatorname{std}(1634)=1423.$
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. Then the left words $\{u_{\mathcal{L}},v_{\mathcal{L}}\}$ form a bipartition of $w_{\mathcal{L}}$ and the right words $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ form a bipartition of $w_{\mathcal{R}}$ and
(2) $d_{R}(w)$ $=d_{R}(w_{\mathcal{L}})+d_{R}(w_{\mathcal{R}}),$
(3) $d_{R}(u)$ $=d_{R}(u_{\mathcal{L}})+d_{R}(u_{\mathcal{R}}),$
(4) $d_{R}(v)$ $=d_{R}(v_{\mathcal{L}})+d_{R}(v_{\mathcal{R}}).$
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$ such that $n\geq 2$ and $w_{n}\neq n$. Then the right words $u_{\mathcal{R}}$ and $v_{\mathcal{R}}$ satisfy the following equation:
(5) $d_{R}(\tilde{u_{\mathcal{R}}})+d_{R}(\tilde{v_{\mathcal{R}}})=d_{R}(u_{\mathcal{R}})+d_{R}(v_{\mathcal{R}})-1.$
@@
The proof is by induction on $\mathbb{N}^{*}$. The result is obviously true in $\mathcal{S}_{1}$ that contains only the identity with 0 descents and inversions. Now, let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$ with $n\geq 2$. We suppose that the statement of Theorem 1.5 is satisfied for all $1\leq k<n$. Write $w=w_{1}\dots w_{n}$ in one-line notation.
If $w_{n}=n$, then $\operatorname{invs}(w)$ does not contain any inversion $(*,n)$. It is also true for $\operatorname{invs}(u)$ and $\operatorname{invs}(v)$ since $\{u,v\}$ is a bipartition of $n$. Therefore $u_{n}=v_{n}=n$. In particular, there is no descent in position $n-1$ for neither $u$, $v$, nor $w$. So $w$, $u$ and $v$ are in $\mathcal{S}_{n-1}$ viewed as a subgroup of $\mathcal{S}_{n}$. By Lemma 3.8, $\{\tilde{u},\tilde{v}\}$ is a bipartition of $\tilde{w}$. So by induction $d_{R}(w)=d_{R}(\tilde{w})=d_{R}(\tilde{u})+d_{R}(\tilde{v})=d_{R}(u)+d_{R}(v)$.
@@
This implies that $w_{\mathcal{L}}$, $u_{\mathcal{L}}$ and $v_{\mathcal{L}}$ are not empty words. Therefore, by Proposition 3.11, $\{u_{\mathcal{L}},v_{\mathcal{L}}\}$ is a bipartition of $w_{\mathcal{L}}$ with $w_{\mathcal{L}}\in\mathcal{S}_{k-1}$, and $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ is a bipartition of $w_{\mathcal{R}}$ with $w_{\mathcal{R}}\in\mathcal{S}_{n-k+1}$. We have $k-1<n$ and $n-k+1<n$, since $1<k<n$,. By induction, we obtain $d_{R}(w_{\mathcal{L}})=d_{R}(u_{\mathcal{L}})+d_{R}(v_{\mathcal{L}})$ and $d_{R}(w_{\mathcal{R}})=d_{R}(u_{\mathcal{R}})+d_{R}(v_{\mathcal{R}})$. Using Eq. (2), Eq. (3), and Eq. (4) of Proposition 3.11, we obtain that $d_{R}(w)=d_{R}(u)+d_{R}(v)$.
@@
This implies that the left words are empty, i.e., $w=w_{\mathcal{R}}$, $u=u_{\mathcal{R}}$, and $v=v_{\mathcal{R}}$ and we cannot use directly the induction hypothesis. Nevertheless, we now have $w_{1}=n$. In particular, as $n\geq 2$, this means that there is a descent in position $1$ in $w$ which disappears if we remove $n$. In other words, $d_{R}(\tilde{w})=d_{R}(w)-1$. Moreover, by Proposition 3.12, we have $d_{R}(\tilde{u})+d_{R}(\tilde{v})=d_{R}(u)+d_{R}(v)-1$. By Lemma 3.8, we know that $\{\tilde{u},\tilde{v}\}$ is a bipartition of $\tilde{w}\in\mathcal{S}_{n-1}$. Therefore, by induction, $d_{R}(\tilde{w})=d_{R}(\tilde{u})+d_{R}(\tilde{v})$. We conclude that:
$d_{R}(w)=d_{R}(\tilde{w})+1=d_{R}(\tilde{u})+d_{R}(\tilde{v})+1=d_{R}(u)+d_{R}(v).$
@@
The definition of the left, right and decreased words of $w$ by choosing $n$ as the letter is crucial in the above proof. A more ‘Coxeter-like’ strategy to prove Theorem 1.5 by induction we tried (and failed) is the following: If $D_{R}(w)=S$, then $w=w_{\circ}$ is the longest element and we can therefore conclude by Proposition 2.5. Now assume that $\tau_{i}\in S\setminus D_{R}(w)$ and define the left, right and decreased words accordingly to $w_{i+1}$. Unfortunately, then the analog of Proposition 3.11 fails to be true.
For instance, consider $w=54231$ with the bipartition $\{u=23514,v=41523\}$. Take $\tau_{3}\in S\setminus D_{R}(w)$, then $w_{4}=3$, $\mathcal{L}(w)=\{542\}$ and $\mathcal{R}(w)=\{1,3\}$. With this definition depending of $3$ instead of $n=5$, we obtain $u_{\mathcal{L}}=132$, $u_{\mathcal{R}}=21$, $v_{\mathcal{L}}=312$ and $v_{R}=12$. Observe that $d_{R}(u)=1<2=d_{R}(u_{\mathcal{L}})+d_{R}(u_{\mathcal{R}})$, which means that Proposition 3.11 fails to be true in this case.
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. Then there is no inversion $(a,c)$ of $u$ or $v$ such that $a\in\mathcal{L}(w)$ is left letter of $w$ and $c\in\mathcal{R}(w)$ is a right letter of $w$. In particular:
  • If $a\in\mathcal{R}(w)$ and $c\in\mathcal{L}(w)$ such that

### v3

@@ 2.3. Bipartitions of the longest element
Recall that $W$ is finite if and only if there is an element of maximal length, which is unique and denoted by $w_{\circ}$. It is well-known that the longest element has the following properties: $\Phi^{+}=\Phi(w_{\circ})$; $w_{\circ}Sw_{\circ}=S$, $D_{R}(w_{\circ})=D_{L}(w_{\circ})=S$. Moreover, for all $w\in W$, we have:
@@
The next statement shows, in particular, that in the case of finite Weyl groups Conjecture 1 is equivalent to Ressayre’s statement $(\triangle)$ from the introduction.
@@ 2.5. Conjecture 1, bipartitions and rectangled intervals in the weak order
The following statement shows that the maximal distance in an interval $e,w]_{R}$ in $(W,\leq_{R})$ is achieved by a geodesic (a smallest path in the right Cayley graph) between $e$ and $w$.
@@
We believe that Proposition 2.12 is well-known, but we could not find a reference so we included a proof. More generally, it is known that the interval $e,w]_{R}$ is convex, i.e., any geodesic between two elements of $e,w]_{R}$ is contained in $e,w]_{R}$. This follows from the interpretation of the inversion set $T(g)$, for $g\in W$, as the set $\mathcal{H}(g)$ of hyperplanes separating $g$ from the identity $e$. One can show that, for $u,v\in[e,w]_{R}$, the set of hyperplanes $\mathcal{H}(u,v)$ that separates $u$ from $v$ satisfies $d(u,v)=|\mathcal{H}(u,v)|$ and $\mathcal{H}(u,v)\subseteq\mathcal{H}(u)\cup\mathcal{H}(v)\subseteq\mathcal{H}(w)$. Moreover, any $g\in W$ in a geodesic from $u$ to $v$ satisfies $\mathcal{H}(u,g)\sqcup\mathcal{(}g,v)=\mathcal{H}(u,v)\subseteq\mathcal{H}(w)$ and $\mathcal{H}(g)\setminus\mathcal{H}(u,v)=H(u)\cup\mathcal{H}(v)\setminus\mathcal{H}(u,v)$. Therefore $\mathcal{H}(g)\subseteq\mathcal{H}(w)$, implying $g\leq_{R}w$.
Proof of Propostion 1.7.
Let $u,v,w\in W$. We need to show that $\{u,v\}$ is a bipartition of $w$ if and only if $\{u,v\}$ is a diameter of $e,w]_{R}$.
Assume first that $\{u,v\}$ is a bipartition of $w$. Therefore $\ell(w)=\ell(u)+\ell(v)$ by Proposition 2.1 (2). Now, by Lemma 2.9, $\Phi(u)\cap\Phi(v)=\emptyset$. Finally we obtain $d(u,v)=\ell(u)+\ell(v)=\ell(w)$ and, therefore, $\{u,v\}$ is a diameter of $e,w]_{R}$.
Assume now that $\{u,v\}$ is a diameter of $e,w]_{R}$. On one hand, by definition and Lemma 2.11, we have:
@@
Together with $(\star)$, we obtain $|\Phi(u)\cap\Phi(v)|\geq 0$. Therefore $\Phi(u)\cap\Phi(v)=\emptyset$. So by $(\star)$ again, we have $\ell(w)=|\Phi(w)|=|\Phi(u)\sqcup\Phi(v)|$. Since $\Phi(u)\cup\Phi(v)\subseteq\Phi(w)$, we conclude that $\Phi(w)=\Phi(u)\sqcup\Phi(v)$. In other words, $\{u,v\}$ is a bipartition of $w$. ∎
We end this section with the following proposition that we need to prove the equivalence between Conjecture 1, Conjecture 2 and Conjecture 3.
@@
In particular, $\{u,v\}$ is a diameter of $e,w]_{R}$ if and only if $\{(\varphi_{w}(u),\varphi_{w}(v)\}$ is a diameter of $e,w^{-1}]_{R}$.
@@
Proof of Theorem 1.8.
The equivalence between Conjecture 1 and Conjecture 2 follows from Proposition 1.7. By Proposition 2.14, the left multiplication by $w^{-1}$ is an order-reversing isomorphism from $e,w]_{R}$ to $e,w^{-1}]_{R}$, i.e., $u\leq_{R}v$ if and only if $w^{-1}v\leq_{R}w^{-1}u$, for all $u,v\in[e,w]_{R}$. In particular, $\operatorname{coatom}([e,w]_{R})=\operatorname{atom}([e,w^{-1}]_{R}$ and, for any $u\leq_{R}w$, $\operatorname{coatom}([e,u]_{R})=\operatorname{atom}([w^{-1}u,w^{-1}]_{R}$. So Conjecture 2 is equivalent to Conjecture 3. ∎
@@ 3. The case of symmetric groups (Type A)
In the two next sections, we provide a direct proof of Conjecture 1 for Coxeter systems of type $A$ (symmetric groups) and $B$ (hyperoctahedral groups). For $n\in\mathbb{N}^{*}$, we denote by $\mathcal{S}_{n}$ the symmetric group of permutations of $\{1,\dots,n\}$ and by $W_{n}$ the hyperoctahedral group of signed permutations of $\bar{n},n]\!]^{*}:=\{\pm 1,\dots,\pm n\}$.
Theorem 3.1.
$d_{R}(w)=d_{R}(u)+d_{R}(v).$
The proof of the type $A$ part of Theorem 3.1 is given in this section and the part regarding type $B$ is given in the next section §4.
Remark 3.2.
Our proof does not require the use of the longest element, we do show directly Conjecture 1 in type $A$ and $B$ by means of shuffle and standardization that seems to suggest that some phenomenon linked to decompositions along minimal coset representatives for spherical standard parabolic cosets might be hidden behind the proof we provide, By providing those proofs ($A$ and $B$) we hope that it might lead to an answer to Conjecture 1 in general.
@@
A permutation $\sigma\in\mathcal{S}_{n}$, as a bijection on the set $n]=\{1,\dots,n\}$, is written as a word on the alphabet $[n]$: $\sigma=\sigma(1)\sigma(2)\dots\sigma(n)$ (the one-line notation). The (left) inversion set of $\sigma$ has therefore the following well-known interpretation:
@@
since in cycle notation (ij)=(ji)(i\ j)=(j\ i). The descent sets are $D_{L}(\sigma)=\{\tau_{3},\tau_{5},\tau_{7}\}$ and $D_{R}(\sigma)=\{\tau_{4},\tau_{5},\tau_{7}\}$. Therefore $\sigma$ has descents in positions $4,5$ and $7$.
@@ 3.1. Standardization of words, inversions and descents
(1) $\operatorname{Inv}(w):=\{(w_{j},w_{i})\mid i<j\text{ and }w_{i}>w_{j}\}=\{(w_{j},w_{i})\mid w=\dots w_{i}\dots w_{j}\dots\}.$
@@
The word $w=391427$ has $2$ descents in positions $2$ and $4$ and $\operatorname{Inv}(w)=\{(1,3),\ (1,9),\ (4,9),\ (2,3),\ (2,9),\ (2,4),\ (7,9)\}$.
@@
In the definition of inversions of the word $w$, we use the notation of a couple $(a,b)$ for an inversion instead of the cycle notation for which (ab)=(ba)(a\ b)=(b\ a). In particular, if $w\in\mathcal{S}_{n}$, then $T(w)=\{(a\ b)\mid(a,b)\in\operatorname{Inv}(w)\}$.
The reason is that, in the proof of Theorem 3.1, we need to keep track of the following fact: $(a,b)\in\operatorname{Inv}(w)$ means that the letter $a$ is strictly smaller that the letter $b$ and the letter $a$ appears after the letter $b$ in $w$, i.e., $w=\dots b\dots a\dots$.
@@
Let $I$ be a set of tuples $(a,b)$ with $1\leq a<b\leq n$ for some $n\geq 1$. Then $I$ is the inversion set of a certain permutation of $\mathcal{S}_{n}$ if and only if it satisfies the following conditions for all $1\leq a<b<c\leq n$
@@
Let $w=526341$. It has $10$ inversions : (1,2),(1,3),(1,4),(1,5),(1,6),(2,5),(3,5),(3,6),(4,5),(4,6)(1,2),(1,3),(1,4),(1,5),(1,6),(2,5),(3,5),(3,6),(4,5),(4,6) and $3$ descents in positions $1$, 33, and $5$. Then $u=234561$ and $v=152634$ form a bipartition of $w$. Indeed, the inversions of $u$ are $(1,2),(1,3),(1,4),(1,5),(1,6)$ while the inversions of $v$ are $(2,5),(3,5),(3,6),(4,5),(4,6)$.
@@
Let $w$ be a permutation of $\mathcal{S}_{n}$ with $n\geq 1$. The decreased word of $w$ is obtained by removing the letter $n$ from $w$ and is denoted by $\widetilde{w}$.
For an example, if $w=526341$, then $\widetilde{w}=52341$.
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$ with $n\geq 1$, then $\{\widetilde{u},\widetilde{v}\}$ is a bipartition of $\widetilde{w}$.
@@
[…]
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. The right letters of $w$ are defined as the set of letters in $w$ at the right-hand side of the letter $n$:
${\overrightarrow{n}}(w):=\{a\mid(a,n)\in\operatorname{Inv}(w)\}\cup\{n\}$
The left letters of $w$ are the complement in $[n]$ of the right letters, i.e., the set of letters in $w$ at the left-hand side of the letter $n$:
${\overleftarrow{n}}(w):=\{b\mid 1\leq b<n,(b,n)\not\in\operatorname{Inv}(w)\}.$
@@
$w_{\overleftarrow{n}}$ $=\operatorname{std}(w_{|{\overleftarrow{n}}(w)}),$ $w_{\overrightarrow{n}}$ $=\operatorname{std}(w_{|{\overrightarrow{n}}(w)})$
$u_{\overleftarrow{n}}$ $=\operatorname{std}(u_{|{\overleftarrow{n}}(w)}),$ $u_{\overrightarrow{n}}$ $=\operatorname{std}(u_{|{\overrightarrow{n}}(w)})$
$v_{\overleftarrow{n}}$ $=\operatorname{std}(v_{|{\overleftarrow{n}}(w)}),$ $v_{\overrightarrow{n}}$ $=\operatorname{std}(v_{|{\overrightarrow{n}}(w)}).$
In other words, we cut $w$ into two parts (left and right) just before the maximal letter. Then we select those letters in $u$ and $v$ and standardize to form new permutations. Note that $n$ is by definition a right letter. The set ${\overleftarrow{n}}(w)$ of left letters might be empty, in this case the left words are the empty word and the right words are the original words $u,v,w$. If $n\geq 1$, the set ${\overrightarrow{n}}(w)$ of right letters is never the empty set as it contains at least the letter $n$. Note also that the left (resp. right) words of $u$ and $v$ depend on the left (resp. right) letters of $w$.
@@
$w_{\overleftarrow{n}}$ $=\operatorname{std}(52)=21,$ $w_{\overrightarrow{n}}$ $=\operatorname{std}(6341)=4231,$
$u_{\overleftarrow{n}}$ $=\operatorname{std}(25)=12,$ $u_{\overrightarrow{n}}$ $=\operatorname{std}(3461)=2341,$
$v_{\overleftarrow{n}}$ $=\operatorname{std}(52)=21,$ $v_{\overrightarrow{n}}$ $=\operatorname{std}(1634)=1423.$
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. Then the left words $\{u_{\overleftarrow{n}},v_{\overleftarrow{n}}\}$ form a bipartition of $w_{\overleftarrow{n}}$ and the right words $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ form a bipartition of $w_{\overrightarrow{n}}$ and
(2) $d_{R}(w)$ $=d_{R}(w_{\overleftarrow{n}})+d_{R}(w_{\overrightarrow{n}}),$
(3) $d_{R}(u)$ $=d_{R}(u_{\overleftarrow{n}})+d_{R}(u_{\overrightarrow{n}}),$
(4) $d_{R}(v)$ $=d_{R}(v_{\overleftarrow{n}})+d_{R}(v_{\overrightarrow{n}}).$
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$ such that $n\geq 2$ and $w_{n}\neq n$. Then the right words $u_{\overrightarrow{n}}$ and $v_{\overrightarrow{n}}$ satisfy the following equation:
(5) $d_{R}(\widetilde{u_{\overrightarrow{n}}})+d_{R}(\widetilde{v_{\overrightarrow{n}}})=d_{R}(u_{\overrightarrow{n}})+d_{R}(v_{\overrightarrow{n}})-1.$
@@
The proof is by induction on $n\in\mathbb{N}^{*}$. The result is obviously true in $\mathcal{S}_{1}$ that contains only the identity with 0 descents and inversions. Now, let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$ with $n\geq 2$. We suppose that the statement of Theorem 3.1 is satisfied for all $1\leq k<n$. Write $w=w_{1}\dots w_{n}$ in one-line notation.
If $w_{n}=n$, then $\operatorname{Inv}(w)$ does not contain any inversion (,n)(*,n). It is also true for $\operatorname{Inv}(u)$ and $\operatorname{Inv}(v)$ since $\{u,v\}$ is a bipartition of $n$. Therefore $u_{n}=v_{n}=n$. In particular, there is no descent in position n1n-1 for neither $u$, $v$, nor $w$. So $w$, $u$ and $v$ are in $\mathcal{S}_{n-1}$ viewed as a subgroup of $\mathcal{S}_{n}$. By Lemma 3.10, $\{\widetilde{u},\widetilde{v}\}$ is a bipartition of $\widetilde{w}$. So by induction $d_{R}(w)=d_{R}(\widetilde{w})=d_{R}(\widetilde{u})+d_{R}(\widetilde{v})=d_{R}(u)+d_{R}(v)$.
@@
This implies that $w_{\overleftarrow{n}}$, $u_{\overleftarrow{n}}$ and $v_{\overleftarrow{n}}$ are not empty words. Therefore, by Proposition 3.13, $\{u_{\overleftarrow{n}},v_{\overleftarrow{n}}\}$ is a bipartition of $w_{\overleftarrow{n}}$ with $w_{\overleftarrow{n}}\in\mathcal{S}_{k-1}$, and $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ is a bipartition of $w_{\overrightarrow{n}}$ with $w_{\overrightarrow{n}}\in\mathcal{S}_{n-k+1}$. We have k1<nk-1<n and nk+1<$n$-k+1<n, since $1<k<n$,. By induction, we obtain $d_{R}(w_{\overleftarrow{n}})=d_{R}(u_{\overleftarrow{n}})+d_{R}(v_{\overleftarrow{n}})$ and $d_{R}(w_{\overrightarrow{n}})=d_{R}(u_{\overrightarrow{n}})+d_{R}(v_{\overrightarrow{n}})$. Using Eq. (2), Eq. (3), and Eq. (4) of Proposition 3.13, we obtain that $d_{R}(w)=d_{R}(u)+d_{R}(v)$.
@@
[…]
$d_{R}(w)=d_{R}(\widetilde{w})+1=d_{R}(\widetilde{u})+d_{R}(\widetilde{v})+1=d_{R}(u)+d_{R}(v).$
@@
The definition of the left, right and decreased words of $w$ by choosing $n$ as the letter is crucial in the above proof. A more ‘Coxeter-like’ strategy to prove Theorem 3.1 by induction we tried (and failed) is the following: If $D_{R}(w)=S$, then $w=w_{\circ}$ is the longest element and we can therefore conclude by Proposition 2.5. Now assume that $\tau_{i}\in S\setminus D_{R}(w)$ and define the left, right and decreased words accordingly to $w_{i+1}$. Unfortunately, then, the analog of Proposition 3.13 fails to be true.
For instance, consider $w=54231$ with the bipartition $\{u=23514,v=41523\}$. Take $\tau_{3}\in S\setminus D_{R}(w)$, then $w_{4}=3$, ${\overleftarrow{n}}(w)=\{542\}$ and ${\overrightarrow{n}}(w)=\{1,3\}$. With this definition depending of $3$ instead of $n=5$, we obtain $u_{\overleftarrow{n}}=132$, $u_{\overrightarrow{n}}=21$, $v_{\overleftarrow{n}}=312$ and $v_{R}=12$. Observe that $d_{R}(u)=1<2=d_{R}(u_{\overleftarrow{n}})+d_{R}(u_{\overrightarrow{n}})$, which means that Proposition 3.13 fails to be true in this case.
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. Then there is no inversion $(a,c)$ of $u$ or $v$ such that $a\in{\overleftarrow{n}}(w)$ is a left letter of $w$ and $c\in{\overrightarrow{n}}(w)$ is a right letter of $w$. In particular:
  • If $a\in{\overrightarrow{n}}(w)$ and $c\in{\overleftarrow{n}}(w)$ such that